If x+3,3x-4,2x+4 are in AP.Find the valve of x.
Answer by richwmiller(17219)

x+3,3x-4,2x+4
2x+4-(3x-4)=(3x-4)-(x+3)
x = 5
2x+4=c (3x-4)=b,(x+3)=a,
c = 14, b = 11, a = 8
difference is 3
add 3